5. A 2.0 cm thick brass plate (kt = 105 W/K m) is sealed to a glass sheet (kt = 0.80 W/K m), and both have the same area. The exposed face of the brass plate is at 80°C, while the exposed face of the glass is at 20 C.

How thick is the glass if the glass brass interface is at 65 C?


1
Expert's answer
2020-11-27T10:16:13-0500
"k_1A\frac{T_1-T}{h_1}=k_2A\frac{T-T_2}{h_2}\\105\frac{80-65}{2}=0.8\frac{65-20}{h_2}\\h_2=0.046\ cm"